Santa Klaus Toys just paid a dividend of $2.50 per share. The required return is 10.2 percent and the perpetual dividend growth rate is 3.4 percent. What price should this stock sell for five years from today?
Answer #1

Price of stock 5 years from now should be   $ 61.78

Step-1:Current price of bond
Current price of bond = D0*(1+g)/(Ke-g) Where,
= 2.50*(1+0.034)/(0.1020-0.034) D0 = $       2.50
= $    38.01 g = 3.40%
Ke = 10.20%
Step-2:Price of bond 5 years from now
Price of bond 5 years from now = P0*(1+Ke)^n Where,
= 38.01*(1+0.1020)^5 P0 = $    38.01
= $    61.78 Ke = 10.20%
n = 5